the product of the page numbers of 2 facing pages of a book is 110. Find the page numbers.
Answer by mananth(16949)

if the pages are facing each other they have to be cosequtive pages.
let one page number be x
the other will be x+1
..
x(x+1)=110
x^2+x-110=0
x^2+11x-10x-110=0
x(x+11)-10(x+11)=0
(x+11)(x-10)=0
x=10 since page number cannot be negative
10 & 11 are the page numbers